For the following reaction, 5.59 grams of oxygen gas are mixed with excess carbon (graphite). The reaction yields 7.10 grams of carbon dioxide.


carbon (graphite) ( s ) + oxygen ( g )  carbon dioxide ( g )


What is the theoretical yield of carbon dioxide? ____ grams

What is the percent yield for this reaction? ____ %


1
Expert's answer
2022-03-08T03:01:03-0500

To determine theoretical yield of "CO_2" ,we need to calculate number of moles of "O_2"


"C_s+O_2\to CO_2"


"\frac{5.59of O_2}{32.0 g/molO_2}=0.175moles of O_2"


"0.175moles of O_2\times\frac{1moleO_2}{1moleCO_2}×44.01g/molof CO_2=7.70175"


"7.70175gof CO_2" is the theoritical yield of "CO_2"


2.The percent yield of "CO_2" is:

Actual yield of CO2 divided by theoretical yield of CO2 ×100%


"=\frac{7.10g}{7.70175}×100\%=92.187\%"